OpenStudy (anonymous):

I will fan or medal, I just really need help please. Solve the system by substitution. -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (parker.goodbar):

x=y+5/4z+−7/4 x=−z+2

OpenStudy (parker.goodbar):

x=−y−z+8

OpenStudy (anonymous):

the answer is (3,6,-1). I just don't know how to get there, and I have to show work... @parker.goodbar

OpenStudy (anonymous):

You are correct :) @imtotallystuck

OpenStudy (anonymous):

Do you know the steps?? @Tyler.Evan2